OpenStudy (anonymous):

Can someone please up I just can not seem to set this problem up right. Solve the triangle. Round lengths to the nearest tenth and angle measure to the nearest degree. a=7, b=7, c=5

OpenStudy (anonymous):

law of cosines for this one

OpenStudy (anonymous):

if you have not got there yet i guess we can solve using some geometry

OpenStudy (anonymous):

so you use a^2 + c^2 - b^2 / 2ac

OpenStudy (anonymous):



OpenStudy (anonymous):

you have the sides, so i take it you are looking for the angles right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you want to use the law of cosines, use \(\cos(\theta)=\frac{a^2+b^2-c^2}{2ab}\)

OpenStudy (anonymous):

I used a^2 + c^2 - b^2 / 2ac and got 5/14 but I don't know what to do now

OpenStudy (anonymous):

take the inverse cosine

OpenStudy (anonymous):

so use cos^-1 on my calulator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so B = 69 degrees

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so two angles are 69 and the other is whatever \(180-2\times 69\) is

OpenStudy (anonymous):

okay thank you I can figure the rest out i was using cos instead of the inverse

OpenStudy (anonymous):

actually \(69.08\) if you want to be picky

OpenStudy (anonymous):

yw